OpenStudy (anonymous):

Two boys want to use a seesaw and they need to move the seesaw so that their weights will balance out. The formula is given by w1 • d1 = w2 • d2 where w1 = weight of the first boy, d1 = distance of the first boy from the fulcrum, w2 = weight of the second boy, and d2 = distance of the second boy from the fulcrum. Rewrite the formula to solve for d2. Show all steps in your work.

OpenStudy (anonymous):

w1 • d1 = w2 • d2 (w1 • d1)/w2 = d2 simply divide by w2 and you have solved for d2

OpenStudy (anonymous):

\[w_{1}d_{1}=w_{2}d_{2}\] Solve for \[d_{2}\] by dividing both sides by \[w_{3}\] Like so: \[\frac{ w_{1}d_{1} }{ w_{2} } = \frac{ w_{2}d_{2} }{ w_{2} }\] As you can see, the w2's cancel out on the right side and now you get \[d_{2} = \frac{ d_{1}w_{1} }{ w_{2} }\] Hope I helped!

OpenStudy (jjuden):

w₁d₁ = w₂d₂ divide both sides by w₂ w₁d₁ / w₂ = d₂